A plate of length 10m width 5m is placed 20m below the surface of sea water of density 1.02g.cm2.if atmospheric pressure is 76cm of mercury, calculate the total force experienced by plate
1
Expert's answer
2021-08-26T14:48:13-0400

"A=10\cdot 5=50\ (m^2)"


"p=\rho gh+101325=1020\cdot9.81\cdot 20+101325=301449\ (Pa)"


"F=p \cdot A=301449\cdot 50=15072450\ (N)\approx15.1\ (MN)"